OpenStudy (anonymous):

How do you set up an equation to find a parabola? I have two given points but I don't know what I'm supposed to be doing? Help Please?

OpenStudy (anonymous):

@phi could you help please?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

hero (hero):

What are the two given points?

hero (hero):

@huntingman15

OpenStudy (anonymous):

(3,0) and (-3,0)

hero (hero):

Those are the zeroes of the parabola. In other words: x = 3 x = -3 Place everything on the same side to get: x - 3 = 0 x + 3 = 0 Multiply (x - 3)(x + 3) to get the parabola expression.

OpenStudy (anonymous):

Is that all I have to do?

hero (hero):

Make sure it is in the form y = f(x) when you are finished.

OpenStudy (anonymous):

Im sorry. Ive been working all day on homework so I a little brain dead. If you multiplied (x-3) (x+3) you would get x squared + -9?

hero (hero):

\((x + 3)(x - 3) = x^2 - 9\) Recall difference of squares rule: \((a + b)(a - b) = a^2 - b^2\)

OpenStudy (anonymous):

I hate to keep bothering you but what would be my next step be then? I don't understand what the difference of squares rule is for?

hero (hero):

It's so that you can multiply \((x + 3)(x - 3)\) in one step. You express the final result as \(y = x^2 - 9\)

OpenStudy (anonymous):

Wow. Okay. I understand now. Thank you so much!

hero (hero):

Use the tagging feature if I don't respond promptly @Hero

hero (hero):

And you're welcome